OpenStudy (anonymous):

What is the answer to 2(x-3)=2x+5

OpenStudy (anonymous):

what??

OpenStudy (debbieg):

Did you try to solve it? what did you do?

OpenStudy (anonymous):

I dont know the answer

OpenStudy (anonymous):

2(x - 3) = 2x+5 2x - 6 = 2x + 5 - 5 -5 2x - 6 - 5 = 2x 2x -11 = 2x can you do the rest?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

Thank

OpenStudy (anonymous):

np

OpenStudy (anonymous):

Hold up am confuse hepl me out

OpenStudy (anonymous):

do you know how to solve for x where the exauation would be for examle 45 = x?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

just need to bring 2 to the other side by deviding 2

OpenStudy (anonymous):

now finish off you should know this

OpenStudy (anonymous):

Ok

OpenStudy (jack1):

@shinebrightlikeadimon ??? this questions doesn't work hey @patrice101 there is no solution to this equation, as per below: 2(x-3) =2x+5 (multiply out the brackets) = 2x - 6 = 2x + 5 (add like terms by moving all the x's to one side) = 2x - 6 - 2x = 2x + 5 - 2x = -6 = + 5 and -6 does not equal 5 ... so ur whole equation doesn't work

OpenStudy (debbieg):

anytime you try to solve and get a result like this one - an equation that no longer involves variables, but only numbers, and is clearly false (-6 = 5), then what you have is an "inconsistent equation" and so it has NO solutions. In other words, there are NO values that can be substituted in for x which will make the equation a true statement. it's easy to see why.... once you get to the point: 2x - 6 = 2x + 5 then add the 6 to both sides 2x = 2x + 11 In this form, it's a bit more intuitively obvious that there can be no solution - no value of x can make 2x = 2x + 11. On the other hand, if you try to solve an equation and reach a point where the variables go away but the equation is TRUE, then the equation is an IDENTITY. E.g.: 2x - 6 = 2(x - 3) 2x - 6 = 2x - 6 subtract 2x from both sides - 6 = -6 add 6 to both sides 0 = 0 Both - 6= -6 and 0 = 0 are TRUE, so you can see that this is an identity. That means that it has INFINITELY MANY solutions, because it is true for EVERY value of x.